if cold water is heated andthe hot water shows pH little less than 7. this is because of (A) hot water become little acidic.
(B) dissociation of water increases
(C) thermal energy is acidic in nature
(D) silicate ions of glass get dissolved
1
Expert's answer
2014-06-06T09:25:57-0400
The dissociation constant of water at ambient temperature is equal to Kw = 1*10-14. When the temperature raises to 100 oC, the Kw value raises by 51 times:
Kw = [H+][OH-] = 51*10-14
[H+] = [OH-], => Kw = [H+]2
[H+] = sqrt(Kw) = 7.16*10-7 mol/l
pH = -lg [H+] = 6.14
We can see that the increase in temperature causes lowering in pH value due to rise in water autoionization constant.

The correct option is (B).